OpenStudy (anonymous):

I need help with two problems I'm stuck on: Solve: -2x - 13 = 8x + 7 -2 2 -6 6 and Solve: -4(2x - 1) = -6(x + 2) - 2 -5 5 9 -9

OpenStudy (anonymous):

for the first problem you add 13 from both sides then subtract 8x from both sides then divide by 10

OpenStudy (anonymous):

the second one im not sure how

OpenStudy (anonymous):

For the second one, you distribute the -4 and -6

OpenStudy (anonymous):

so 13+7=20 -2x=8x+20 ?

OpenStudy (anonymous):

You multiply 2x by -4 giving you -8x, and -1 by -4 giving you 4. (for the left side of the equation.)

OpenStudy (anonymous):

yes then subtract 8x on both sides

OpenStudy (anonymous):

what would you get when you subtract 8x from -2x

OpenStudy (anonymous):

so -2x-8x= -10

OpenStudy (anonymous):

or positive 10?

OpenStudy (anonymous):

oh so 2!

OpenStudy (anonymous):

briajosephs Im a little confused about the second one

OpenStudy (anonymous):

The answer's negative two for the first equation.

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

no it would be -10x then divide by -10 and what do you get?

OpenStudy (anonymous):

Oh and for the second equation, the hardest part part is to distribute the outer numbers into the parenthesis

OpenStudy (anonymous):

I got 2, but it was negative 2

OpenStudy (anonymous):

there is your answer.x=-2

OpenStudy (anonymous):

Oh okay! I understand the second on know, give me a scond to solve it

OpenStudy (anonymous):

This website explains the distributive property much better than I can lol http://www.coolmath.com/prealgebra/06-properties/05-properties-distributive-01.htm

OpenStudy (anonymous):

do you get how to solve the first equation now?

OpenStudy (anonymous):

yes I understand the first one, Thanks:) For the second on I got -8x- (-4)=(-6x+-12)-2 I don't know if that was the right step

OpenStudy (anonymous):

It was! Two negatives equal a positive, so it'll be -8x + 4

OpenStudy (anonymous):

i would just go to the website briajosephs gave cause i dont undertand the second equation that well

OpenStudy (anonymous):

oh okay! Then I just solve what's in parentheses right?

OpenStudy (anonymous):

Your new equation is -8x + 4 + -6x - 12 - 2 really because to solve the parenthesis you'd have to distribute again and there's really no need for that

OpenStudy (anonymous):

-8x + 4 = -6x - 12 - 2* sorry lol

OpenStudy (anonymous):

Now you just combine like terms on the right -8x + 4 = -6x -14 Then subract -8 on both sides 4 = 2x -14 And add 14 on both sides 18 = 2x Then divide two and you got your answer!

OpenStudy (anonymous):

So the answer is 9! Thank you so much! you really helped ^-^

OpenStudy (anonymous):

Aw no problem love! Any time (: